OpenStudy (anonymous):

cal culate the frequency of the note produced by 100 cm wire when streched by a load of 100kg if 2cm of wire has a mass of 12.25g....help meeeeee!!!! the answer is 20hz

OpenStudy (shikamaru11):

well first we've to find out the mass of 100cm wire....... if 2cm wire.........=12.25g then 1cm...........=12.25/2 so 100cm.............=12.25/2 x100=612.5g now we'are provided with 100kg load so F=mg F=100x10=1000N and \[v=\sqrt{F\m}\] so we can find the velocity and then by using relation v=fl Or f=v/l................ but what's the note 1st 2nd or 3rd????or ................. etc then you can sort it out easily you can also use \[f=1/2l*\sqrt{F\m}\] ........... hope it'll help :\

OpenStudy (irishboy123):

this question involves the use of the equation for the speed (v) of propagation of a wave in a string that is under tension (T) and with **linear** density \(\mu \), which is : \(\huge v = \sqrt{T \over \mu}\) "2cm of wire has a mass of 12.25g" --> calculate \(\mu\), careful with units "streched by a load of 100kg " --> T = 100g the formula gives you v if i were you i would then use the first harmonic:  from \( {\lambda \over 2 }= 1m\) and \(v = f \lambda, \; f = { v \over \lambda}\) , you should get a frequency it's all here together with derivation: https://en.wikipedia.org/wiki/Vibrating_string